OpenStudy (wintersuntime):

Can someone be nice enough to help me with these worksheet front and back

jimthompson5910 (jim_thompson5910):

If it's possible, please use your phone or some kind of camera to take a snapshot of the worksheets. Then post the images of the worksheets to openstudy. Thank you

OpenStudy (wintersuntime):

It's taking forever

jimthompson5910 (jim_thompson5910):

that's ok. Sometimes tech takes a bit of time to load

OpenStudy (wintersuntime):

Taking forever !!!

OpenStudy (wintersuntime):

Thats the front page I need help with all of them

OpenStudy (wintersuntime):

The question says 1. For each system below, decide which strategy to use. That is, which method would be the most efficient, convenient, and accurate?

OpenStudy (wintersuntime):

@PillowFace750

jimthompson5910 (jim_thompson5910):

ok one moment while I look it over

OpenStudy (wintersuntime):

okay

jimthompson5910 (jim_thompson5910):

ok seems reasonable. Did you have an idea how to start or are you completely stumped on the first step?

OpenStudy (wintersuntime):

I'm confused

jimthompson5910 (jim_thompson5910):

which is 1a closely similar to compared to the examples in the box above?

OpenStudy (wintersuntime):

Substitution ?

jimthompson5910 (jim_thompson5910):

yep, you are correct

jimthompson5910 (jim_thompson5910):

start with the second equation `3x-2y = 4` then replace all of `x` with `(4-2y)` which is the substitution step

jimthompson5910 (jim_thompson5910):

so `3x-2y = 4` turns into `3(4-2y)-2y = 4` Do you see how to solve for y?

OpenStudy (wintersuntime):

Then distribute

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (wintersuntime):

I got 12-6y-2y= 4

jimthompson5910 (jim_thompson5910):

good

OpenStudy (wintersuntime):

Then there are two like terms so I have to add them

jimthompson5910 (jim_thompson5910):

which gets you what?

OpenStudy (wintersuntime):

8

jimthompson5910 (jim_thompson5910):

not just 8, but you're close in a way

OpenStudy (wintersuntime):

12-8y=4 ?

jimthompson5910 (jim_thompson5910):

much better

jimthompson5910 (jim_thompson5910):

then what?

OpenStudy (wintersuntime):

add 8 on both sides ??

OpenStudy (wintersuntime):

Add 12 on both sides ??

jimthompson5910 (jim_thompson5910):

not add, but you're close

OpenStudy (wintersuntime):

subtract

jimthompson5910 (jim_thompson5910):

if it helps, think of 12-8y as -8y+12

jimthompson5910 (jim_thompson5910):

yes subtract 12 from both sides

OpenStudy (wintersuntime):

y=1?

OpenStudy (wintersuntime):

and x=2 ?

jimthompson5910 (jim_thompson5910):

yep y = 1

jimthompson5910 (jim_thompson5910):

and x = 2, good

OpenStudy (wintersuntime):

Now I need help with the second one

jimthompson5910 (jim_thompson5910):

ok one second

jimthompson5910 (jim_thompson5910):

hopefully you agree that 1b matches with the elimination form?

OpenStudy (wintersuntime):

So its elimination right

jimthompson5910 (jim_thompson5910):

yes, if you were to add the equations up, notice how the y terms combine to y+y = 2y we want them to combine to 0y or 0. What change can we make so that the y terms add to 0?

OpenStudy (wintersuntime):

subtract

jimthompson5910 (jim_thompson5910):

or put another way, multiply one of the terms by -1, then add, agreed?

OpenStudy (wintersuntime):

okay

jimthompson5910 (jim_thompson5910):

ultimately it doesn't matter how you think of it because the two are equivalent

OpenStudy (wintersuntime):

So 1b its elimination with multiplication ??

jimthompson5910 (jim_thompson5910):

yes because we cannot add first until we multiply one of the equations (both sides) by -1

OpenStudy (wintersuntime):

So it looks like this y+y=2y)-1

jimthompson5910 (jim_thompson5910):